### v1

@@ Abstract
We explored two parameter-efficient fine-tuning techniques (LoRa and IA3) on two pre-trained LLMs for code generation. We crawled a fine-tuning dataset (14,622 C/C++ files) for secure code generation by collecting code fixes of confirmed vulnerabilities from open-source repositories. Our evaluation dataset comprises 52 vulnerability scenarios designed to cover the top most dangerous C/C++ Common Weakness Enumerations (CWEs). Each scenario is a prompt that may induce LLMs to generate vulnerable code. Our exploration reveals that fine-tuning LLMs can improve secure code generation by 6.4% in C language and 5.4% in C++ language. We further experimented with fine-tuning LLMs using different versions of the collected secure code dataset (block, function, and line). We found that fine-tuning with function-level and block-level datasets achieves the best secure code generation performance, compared to the alternatives (file-level and line-level). In addition, we investigate the impact of the size of the fine-tuning dataset on secure code generation. With the increasing size of the fine-tuning dataset, the model can produce more valid and secure code. Finally, our experiment shows that the LLMs fine-tuned for secure code generation do not have performance degradation in terms of generating correct code. For CodeLlama, we even observed an improvement of 2% in terms of the PASS@1 under the HumanEval CPP benchmark with secure fine-tuning.
Keywords:
Code Generation Cybersecurity Artificial Intelligence (AI) Common Weakness Enumerations (CWEs) Large Language Models (LLMs)
@@ 1 Introduction
Large Language Models (LLMs) trained on huge corpus are significantly advancing various software engineering tasks, including code and comment generation Ahmed and Devanbu (2022); Luo et al. (2023), program repair Jin et al. (2023); Joshi et al. (2023), and bug detection Wu et al. (2024). Specifically in code generation, GPT-4 Turbo has achieved a score of 85.4% on the pass@1 metric of the widely recognized HumanEval Chen et al. (2021) code generation benchmark. Similarly, a recent open-source LLM, MagicoderS-CL, has scored a pass@1 of 70.7% on the same benchmark Wei et al. (2023).
@@
There have been several attempts to promote the generation of secure code by LLMs. Classic vulnerability detection and repair techniques can be applied to detect and fix vulnerabilities CodeQL ([n. d.]). An alternative direction is to improve LLMs through fine-tuning techniques. He et al. He and Vechev (2023) propose a fine-tuning technique based on prefix tuning for LLMs that aims to reduce the generation of vulnerable code. This work is limited to one LLM, i.e., CodeGen Nijkamp et al. (2023b) with 2.7B parameters, and considers specific vulnerabilities (only 9 CWEs) for evaluation. Our prior work Li et al. (2024) fine-tuned GPT-J using the 4,900 vulnerability-fixing commits. However, the full potential of fine-tuned LLMs for secure code generation has not yet been extensively explored. In particular, our previous study was limited in scope, focusing only on a single model and dataset. Therefore, there is a need for further research to (1) experiment with a variety of fine-tuning techniques, (2) experiment with diverse LLMs, (3) investigate the impact of fine-tuning datasets in terms of granularity and size on the performance of fine-tuned LLMs, and (4) validate whether the fine-tuned LLMs for better security still have comparable performance in terms of functional correctness. In other words, fine-tuning LLMs for security purposes does not compromise their ability to generate functionally correct code. To address these gaps, the work aim to provide a comprehensive understanding of whether fine-tuning LLMs would lead to secure code generation in this work. We extend our previous reserach by experimenting with two LLMs (CodeGen2 and CodeLlama) for code generation and leveraging two fine-tuning techniques, i.e., LoRA Hu et al. (2021) and IA3 Liu et al. (2022). Additionally, we try different granularities of the fine-tuning sets. They are file-level, function-level, and block-level, and line-level.
@@
To achieve our goal, we followed a five-step process. First, we curated a dataset of 4,900 vulnerability-fixing commits from 580 open-source projects, sourcing from three widely-used vulnerability datasets (Bhandari et al., 2021; Challande et al., 2022; Fan et al., 2020) as our fine-tuning dataset. Second, we chose CodeLlama (Roziere et al., 2023), and CodeGen2 (Nijkamp et al., 2023a) as the base LLMs for our study and fine-tuned these models using two parameter-efficient fine-tuning (PEFT) techniques, i.e., LoRA and IA3. Third, to evaluate the LLMs’ capability in generating secure code, we created 29 security-sensitive scenarios, in addition to reusing the 23 scenario dataset proposed by Pearce et al. Pearce et al. (2022). In total, there are 52 security-sensitive scenarios. These scenarios cover the top 11 CWE (MITRE’s Common Weakness Enumerations) from the “2022 CWE Top 25 Most Dangerous Software Weaknesses” list  [^3] https://cwe.mitre.org/top25/archive/2022/2022_cwe_top25.html. Fourth, we utilized a static vulnerability detection tool, CodeQL (CodeQL, [n. d.]), to identify vulnerabilities in the generated code by LLMs. After getting the vulnerability results, we conducted a comparative evaluation between the pre-trained and their fine-tuned versions. Furthermore, we assessed the impact of different fine-tuning datasets, specifically focusing on granularity and size, on the performance of the fine-tuned LLMs in generating secure code. Finally, we evaluated the functional correctness of code generated by the pre-trained and fine-tuned models using the HumanEval_CPP benchmark (Zheng et al., 2023). Our exploratory study addresses the following four research questions (RQs):
@@
  • RQ1: Can fine-tuning LLMs with secure source code reduce vulnerable code by LLMs? We leveraged two popular PEFT techniques, LoRA and IA3 to fine-tune CodeLlama and CodeGen2. Our experiment reveals that fine-tuned LLMs have an improved secure ratio for code generation in C and C++. For example, CodeGen2 (fine-funed by LoRa) improves the percentage of secure code from 56% to 62.4%.
@@
  • RQ2: How the granularity of fine-tuning dataset may impact fine-tuning LLMs for secure code generation? For each LLM and PEFT technique, we experimented with four fine-tuning datasets with diverse granularity, i.e., file-level, method-level, block-level, and line-level, to fine-tuning the model. We found that by using block-level and function-level fine-tuning datasets the two models achieve the highest secure ratio with 60.7% and 65.1% for C and C++ in the fine-tuned CodeLlama.
  • RQ3: How the size of fine-tuning dataset may impact fine-tuning LLMs for secure code generation? Similar to RQ2, we evaluated whether fine-tuning datasets of different sizes will yield fine-tuned LLMs with different performances in secure code generation. We find that, by increasing the size of the dataset, the models can generate more secure code snippets from 62.5% to 79.2% in C using LoRa for CodeGen2.
@@
  • RQ4: Does fine-tuning LLMs for security impact the performance of LLMs in generating correct code? We evaluated the performance of the fine-tuned LLMs in generating correct code using the HumanEval_CPP dataset Zheng et al. (2023). We found that the fine-tuned LLMs for secure code generation do not have significant performance degradation in generating correct code. Several of the fine-tuned LLMs from CodeGen2 for C++ even show improvements on PASS@1.
@@
    • We experimented with a variety of dataset versions (in terms of granularity and size) to fine-tune LLMs and evaluate the impacts on generating secure and functionally correct code. We found that using LoRa to fine-tune block-level and function-level data sets achieves the highest performance in C and C++ languages, with 6.7% and 5.4% improvements, respectively. Besides, the functional correctness of our fine-tuned models does not have a significant decrease compared to the pre-trained ones.
    • We release a replication package [^4] https://github.com/SecureLLM/Secure_LLM
    The rest of the paper is organized as follows: Section 2 presents the related work. Section 3 presents in detail our selected subjects (the LLMs and the datasets selected for fine-tuning, and the prompt dataset for evaluation) in this work. In Section 4, we overview our experiment design and describe the detailed configurations/settings we used for fine-tuning LLMs and using the fine-tuned LLMs for secure code generation. Section 5 presents the motivation, approach, and results of three research questions. Section 6 discusses the internal and external threats in this study. Finally, Section 7 concludes the entire study with findings.
@@ 2.1 Large Language Models (LLMs) for Code Generation
    There has been increasing interest in using pre-trained LLMs for code generation and completion. So far, three types of LLMs have been adopted to generate and complete code automatically, namely decoding-only language models, masked language models (MLMs), and encoder-decoder models. One of the most popular examples in decoding-only models is GPT and its series (GPT2, GPT-3, GPT-3.5, and GPT-4) by OpenAI. Besides, GPT-J (6B) Wang and Komatsuzaki (2021), Codex (12B) Chen et al. (2021), CodeGen Nijkamp et al. (2023b, a) and CodeLLaMA Roziere et al. (2023) are decoding-only models, which generate the code according to the previously given tokens. CuBERT (345B) Kanade et al. (2020), CodeBERT (125M) Feng et al. (2020), and InCoder Fried et al. (2022) are MLMs, which target on predicting the masked code according to the code context around the masked pieces. CodeT5 (220M) Wang et al. (2021) is an encoder-decoder model, where an input sequence was read in entirety and encoded to an internal representation and decoded to generate an output sequence leveraging a decoding-only model.
@@ 2.2 Security Issue in LLM-generated Code
    Although LLMs offer significant potential for automating code generation, they are mostly trained on unsanitized data from open-source repositories hosted on GitHub. This practice neglects explicit considerations for security, thereby increasing the risk of propagating existing security vulnerabilities through the generated code. Researchers have highlighted the security risks associated with code generated by LLMs. For instance, Pearce et al. Pearce et al. (2022) evaluated GitHub Copilot, which is based on the OpenAI Codex model, and reported that approximately 40% of the generated code was insecure. Khoury et al. Khoury et al. (2023) analyzed security-relevant coding scenarios with ChatGPT, finding that it produced insecure code in 16 cases, of which only 7 were corrected after further interaction with the model. Moreover, Perry et al. Perry et al. (2023) indicated that developers assisted by AI models tend to introduce more security vulnerabilities than those who do not use such tools.
@@ 2.4 Datasets of Vulnerabilities and Fixes
    Vulnerability databases play a crucial role in analyzing vulnerabilities and evaluating automated approaches (e.g., vulnerability detection and repair) as they collect, maintain, and disseminate information on security vulnerabilities. The National Vulnerability Database (NVD) NVD (2024) is the most comprehensive publicly accessible repository of vulnerabilities. NDV is built upon Common Vulnerability and Exposures (CVE) CVE-MIRTE (2024) entries, which provide a unique identifier for each security flaw, facilitating the sharing of information across different entities. Many studies leverage NVD reports and CVE entries to create datasets for data-driven approaches to vulnerability detection and prediction Sun et al. (2021). Additionally, the Common Weakness Enumeration (CWE) CWE-MIRTE (2022), maintained by MITRE, categorizes types of weaknesses in software and hardware, using a tree-like architecture to link related vulnerabilities. In NDV, the identified CVE of one project can be linked to CWE to indicate the category of weakness.
@@ 3 Study Subjects and Datasets
    In this section, we describe the study subjects and datasets in detail, including two LLMs for code generation (i.e., CodeLlama, and CodeGen2), a collection of real-world secure code (i.e., vulnerability fixes) as the fine-tuning dataset, and an evaluation dataset for evaluating the security aspects of LLM-based code generation.
@@ 3.1 Selected Large Language Models
    Among the many LLMs Wang et al. (2021); Xu et al. (2022) available for code generation, we select two representative medium-size LLMs with around 7 billion parameters: CodeGen2-7BNijkamp et al. (2023a), and CodeLlama-7BRoziere et al. (2023). These models are designed for code generation and strike a balance between computational efficiency and performance.
    CodeGen2 Nijkamp et al. (2023a) is a family of autoregressive language models that focus on the program synthesis, trained on a dataset named “The Stack” (Kocetkov et al., 2022). The authors provide four versions (1B, 3.7B, 7B, and 16B) of CodeGen2. We select the 7B version of CodeGen2 in this study.
    CodeLlama Roziere et al. (2023) has the same architecture as Llama 2 (Touvron et al., 2023) but is further trained specifically on code, infilling code, and human instruction datasets by Meta, aiming at programming tasks. It also supports long input contexts with 100,000 tokens. It provides three versions of parameters (7B, 13B, 34B) with three different variants (i.e., CodeLlama base model for general code generation, CodeLlam-Python for Python code, and CodeLlam-Instruct for safely using in coding assistant). We select the 7B version of CodeLlama in this study.

### v2

@@ Abstract
We explored full fine-tuning and two parameter-efficient fine-tuning techniques (LoRA and IA3) on four pre-trained LLMs for code generation. We crawled a fine-tuning dataset (14,622 C/C++ files) for secure code generation by collecting code fixes of confirmed vulnerabilities from open-source repositories. Our evaluation dataset comprises 52 vulnerability scenarios designed to cover the top most dangerous C/C++ Common Weakness Enumerations (CWEs). Each scenario is a prompt that may induce LLMs to generate vulnerable code. Our exploration reveals that fine-tuning LLMs using PEFT techniques can enhance secure code generation. We observe maximum improvements in security of 6.4% in C language and 5.0% in C++ language. In addition, we compared between the fine-tuning approaches and the prompt-based approaches. The LoRA-tuned models outperform the prompt-based approaches in secure code generation. We further experimented with fine-tuning LLMs using different versions of the collected secure code dataset (file, block, function, and line). We found that fine-tuning with function-level and block-level datasets achieves the best secure code generation performance, compared to the alternatives (file-level and line-level).
Finally, our experiment shows that the LLMs fine-tuned for secure code generation do not have performance degradation in terms of generating correct code. For CodeLlama, we even observed an improvement of 2% in terms of the PASS@1 under the HumanEval_CPP benchmark after fine-tuning.
@@ 1 Introduction
Large Language Models (LLMs) trained on a huge corpus are significantly advancing various software engineering tasks, including code and comment generation Ahmed and Devanbu (2022); Luo et al. (2023a), program repair Jin et al. (2023); Joshi et al. (2023), and bug detection Wu et al. (2024). Specifically in code generation, GPT-4 Turbo has achieved a score of 85.4% on the PASS@1 metric of the widely recognized HumanEval Chen et al. (2021) code generation benchmark. Similarly, a recent open-source LLM, MagicoderS-CL, has scored a PASS@1 of 70.7% on the same benchmark Wei et al. (2023).
@@
There have been several attempts to promote the generation of secure code by LLMs. A direction is to improve LLMs through fine-tuning techniques. For instance, He et al. He and Vechev (2023) proposed a method based on prefix tuning to reduce vulnerable code generation. Another work Li et al. (2024) fine-tuned GPT-J using 4,900 vulnerability-fixing commits. However, these existing studies on fine-tuning LLMs have limited generalization on the (1) subject LLMs, (2) limited exploration on diverse fine-tuning techniques, (3) limited evaluation dataset, and (4) limited in-depth experiments on the impacts of training dataset specifics. First, He et al. He and Vechev (2023) and Li et al. Li et al. (2024) include a single LLM. Second, prior works explore limited fine-tuning techniques, such as full fine-tuning in Li et al. Li et al. (2024) and prefix-tuning in He et al. He and Vechev (2023). Third, the prior work only focuses narrowly on a limited set of vulnerabilities (e.g., only 9 CWE in He et al. He and Vechev (2023). Fourth, the different fine-tuning techniques have not been explored in prior works. Other studies Hajipour et al. (2023); Zhang et al. (2024) rely on prompt engineering rather than modifying the model’s internal weights, limiting their robustness and general applicability. This work systematically addresses these challenges through (1) extending our previous research by experimenting with four LLMs (CodeGen2, CodeLlama, Magicoder, and DeepSeek-Coder) for code generation; (2) exploring fine-tuning methods by leveraging three fine-tuning techniques, i.e., LoRA Hu et al. (2021), IA3 Liu et al. (2022), and full fine-tuning; (3) extending the evaluation dataset to 52 security-related prompts; (4) exploring different granularities of the fine-tuning sets. They are file-level, function-level, block-level, and line-level.
@@
In particular, we followed a five-step process. First, we curated a dataset of 4,900 vulnerability-fixing commits from 580 open-source projects, sourcing from three widely-used vulnerability datasets (Bhandari et al., 2021; Challande et al., 2022; Fan et al., 2020) as our fine-tuning dataset. Second, we chose CodeLlama (Roziere et al., 2023), CodeGen2 (Nijkamp et al., 2023a), DeepSeek-Coder (Guo et al., 2024), and Magicoder (Wei et al., 2023) as the base LLMs for our study and fine-tuned these models using full fine-tuning and two parameter-efficient fine-tuning (PEFT) techniques, i.e., LoRA and IA3. Third, to evaluate the LLMs’ capability in generating secure code, we created 29 security-sensitive scenarios, in addition to reusing the 23 scenario dataset proposed by Pearce et al. Pearce et al. (2022). In total, there are 52 security-sensitive scenarios. These scenarios cover the top 10 CWE (MITRE’s Common Weakness Enumerations) from the “2024 CWE Top 25 Most Dangerous Software Weaknesses” list  [^3] https://cwe.mitre.org/top25/archive/2024/2024_cwe_top25.html. Fourth, we utilized a static vulnerability detection tool, CodeQL (CodeQL, [n. d.]), to identify vulnerabilities in the generated code by LLMs. In our study, we define secure code as code that is free from known vulnerabilities. After getting the vulnerability results, we conducted a comparative evaluation between the pre-trained and their fine-tuned versions. Furthermore, we assessed the impact of different fine-tuning datasets, specifically focusing on granularity and size, on the performance of the fine-tuned LLMs in generating secure code. Finally, we evaluated the functional correctness of code generated by the pre-trained and fine-tuned models using the HumanEval_CPP benchmark (Zheng et al., 2023). Our exploratory study addresses the following four research questions (RQs):
@@
  • RQ1: Can fine-tuning LLMs with secure source code reduce vulnerable code by LLMs? We leveraged full fine-tuning and two popular PEFT techniques, LoRA and IA3 to fine-tune CodeGen2, CodeLlama, Magicoder, and DeepSeek-Coder. Our experiment reveals that fine-tuned LLMs have an improved secure ratio for code generation in C and C++. For example, CodeGen2 (fine-tuned by LoRA) improves the percentage of secure code from 56.0% to 62.4%.
  • RQ2: How does fine-tuning compare with prompt-based methods? We performed a comparative evaluation between the the PEFT techniques and the prompt-based methods. The results shows that the Generic Security and Specific Security prompts achieved average secure ratios of 59.9% and 57.5%, respectively. Both are slightly lower than the secure ratio achieved by the LoRA-tuned approach at 60.9%.
@@
  • RQ3: How the granularity of fine-tuning dataset may impact fine-tuning LLMs for secure code generation? For each LLM and PEFT technique, we experimented with four fine-tuning datasets with diverse granularity, i.e., file-level, method-level, block-level, and line-level, to fine-tuning the model. We found that by using block-level and function-level fine-tuning datasets, the four models improved secure ratios in both C and C++. Specifically, the LoRA-tuned CodeLlama achieved secure ratios of 60.3% for C and 63.1% for C++ on function-level dataset, while the LoRA-tuned DeepSeek-Coder reached 59.5% for C and 63.3% for C++. CodeGen2 attained 60.1% for C and 61.0% for C++ after function-level fine-tuning, and Magicoder achieved 58.7% for C and 66.0% for C++.
@@
  • RQ4: Does fine-tuning LLMs for security impact the performance of LLMs in generating correct code? We evaluated the performance of the fine-tuned LLMs in generating correct code using the HumanEval_CPP dataset Zheng et al. (2023). We found that fine-tuned LLMs for secure code generation do not have significant performance degradation in generating correct code. Interestingly, models such as CodeLlama, Magicoder, and DeepSeek-Coder, when fine-tuned using IA3 with a function-level dataset, outperformed their pretrained models and demonstrated improvements in PASS@1.
@@
    • We experimented with a variety of dataset versions (in terms of granularity and size) to fine-tune LLMs and evaluate the impacts on generating secure and functionally correct code. We found that using LoRA to fine-tune block-level and function-level data sets achieves the highest performance in C and C++ languages, with maximum improvements of 6.4% and 5.9%, respectively. Besides, the functional correctness of our fine-tuned models does not have a significant decrease compared to the pre-trained ones.
    • We release a replication package [^4] https://github.com/SecureLLM/Secure_LLM.
    The rest of the paper is organized as follows: Section 2 presents the related work. Section 3 presents in detail our selected subjects (the LLMs and the datasets selected for fine-tuning, and the prompt dataset for evaluation) in this work. In Section 4, we overview our methodology and describe the detailed design of our experiment. Section 5 presents the motivation, approach, and results of four research questions. Section 6 discusses extensions of our study, including experiments with additional programming languages, qualitative analyses, and the application of PEFT to a larger model. Section 7 discusses the internal and external threats in this study. Section 8 summarizes the insights and the future directions of the study. Finally, Section 9 concludes the entire study.
@@ 2.1 Large Language Models (LLMs) for Code Generation
    There has been increasing interest in using pre-trained LLMs for code generation and completion. So far, three types of LLMs have been adopted to generate and complete code automatically, namely decoding-only language models, masked language models (MLMs), and encoder-decoder models. One of the most popular examples in decoding-only models is GPT and its series (GPT2, GPT-3, GPT-3.5, and GPT-4) by OpenAI. Besides, GPT-J (6B) Wang and Komatsuzaki (2021), Codex (12B) Chen et al. (2021), CodeGen Nijkamp et al. (2023b, a) and CodeLlama Roziere et al. (2023) are decoding-only models, which generate the code according to the previously given tokens. CuBERT (345B) Kanade et al. (2020), CodeBERT (125M) Feng et al. (2020), and InCoder Fried et al. (2022) are MLMs, which aim to predict the masked code according to the code context around the masked pieces. CodeT5 (220M) Wang et al. (2021) is an encoder-decoder model, where an input sequence was read in entirety and encoded to an internal representation and decoded to generate an output sequence leveraging a decoding-only model.
    2.2 Improving and Evaluating LLM Code Generation
    The research community has made efforts on LLM code generation on improvement methods and evaluation beyond PASS@K. On the improvement side, specification-driven prompting sharpens intent alignment for transformations such as code translation, revealing gaps between natural-language specifications and faithful execution Saha et al. (2024). Agentic pipelines that pair contextual retrieval with role-specialized collaboration further strengthen method-level refactoring by reducing reasoning slip-ups and stabilizing edits Xu et al. (2025). On evaluation, robustness is increasingly tested along axes developers care about: non-functional requirements (e.g., performance, memory, security) expose brittleness that functional tests miss Lin et al. (2025), and behavior varies across programming languages, complicating generalization claims and motivating cross-language reporting Rabbi et al. (2025). Ethical and sociotechnical risks also surface, i.e., social bias can emerge in generated names, comments, and policy defaults, arguing for bias-sensitive benchmarks and mitigations in the evaluation loop Ling et al. (2025). Methodologically, white-box testing insights from machine translation suggest coverage-like and sensitivity-guided test selection for code generation systems that expose intermediate plans or signals Shao et al. (2025).
@@ 2.3 Security Issue in LLM-generated Code
    Although LLMs offer significant potential for automating code generation, they are mostly trained on unsanitized data from open-source repositories hosted on GitHub. This practice neglects explicit considerations for security, thereby increasing the risk of propagating existing security vulnerabilities through the generated code. Researchers have highlighted the security risks associated with code generated by LLMs. For instance, Pearce et al. Pearce et al. (2022) evaluated GitHub Copilot, which is based on the OpenAI Codex model, and reported that approximately 40% of the generated code was insecure. Khoury et al. Khoury et al. (2023) analyzed security-relevant coding scenarios with ChatGPT, finding that it produced insecure code in 16 cases, of which only 7 were corrected after further interaction with the model. Moreover, Perry et al. Perry et al. (2023) indicated that developers assisted by AI models tend to introduce more security vulnerabilities than those who do not use such tools. Another study Mastropaolo et al. (2024) explored the impact of different strategies in bug-fixing of LLMs. Results show that supervised training on the bug-fixing dataset can benefit the LLMs over the fine-tuning.
@@ 2.5 Datasets of Vulnerabilities and Fixes
    Vulnerability databases play a crucial role in analyzing vulnerabilities and evaluating automated approaches (e.g., vulnerability detection and repair) as they collect, maintain, and disseminate information on security vulnerabilities. The National Vulnerability Database (NVD) NVD (2024) is the most comprehensive publicly accessible repository of vulnerabilities. NVD is built upon Common Vulnerability and Exposures (CVE) CVE-MITRE (2024) entries, which provide a unique identifier for each security flaw, facilitating the sharing of information across different entities. Many studies leverage NVD reports and CVE entries to create datasets for data-driven approaches to vulnerability detection and prediction Sun et al. (2021). Additionally, the Common Weakness Enumeration (CWE) CWE-MITRE (2022), maintained by MITRE, categorizes types of weaknesses in software and hardware, using a tree-like architecture to link related vulnerabilities. In NVD, the identified CVE of one project can be linked to CWE to indicate the category of weakness.
@@ 3 Study Subjects and Datasets
    In this section, we describe the study subjects and datasets in detail, including four LLMs for code generation (i.e., CodeLlama, CodeGen2, Magicoder, and DeepSeek-Coder), a collection of real-world secure code (i.e., vulnerability fixes) as the fine-tuning dataset, and an evaluation dataset for evaluating the security aspects of LLM-based code generation.
@@ 3.1 Selected Large Language Models
    Among the many LLMs Wang et al. (2021); Xu et al. (2022) available for code generation, we select four representative medium-size LLMs with around 7 billion parameters: CodeGen2-7B Nijkamp et al. (2023a), CodeLlama-7B Roziere et al. (2023), Magicoder-6.7B Wei et al. (2023) and DeepSeek-Coder-7B Guo et al. (2024). These models are designed for code generation and strike a balance between computational efficiency and performance.
    CodeGen2 Nijkamp et al. (2023a) is a family of autoregressive language models that focus on program synthesis, trained on a dataset named “The Stack” (Kocetkov et al., 2022). The authors provide four versions (1B, 3.7B, 7B, and 16B) of CodeGen2. We select the 7B version of CodeGen2 in this study.
    CodeLlama Roziere et al. (2023) has the same architecture as Llama 2 (Touvron et al., 2023) but is further trained specifically on code, infilling code, and human instruction datasets by Meta, aiming at programming tasks. It also supports long input contexts with 100,000 tokens. It provides three versions of parameters (7B, 13B, 34B) with three different variants (i.e., CodeLlama base model for general code generation, CodeLlama-Python for Python code, and CodeLlama-Instruct for safely using in coding assistant). We selected the 7B version of CodeLlama in this study.
    Magicoder Wei et al. (2023) is instruction‑tuned using OSS‑Instruct designed for coding tasks. There are two variants including Magicoder‑7B (based on CodeLlama) and Magicoder‑6.7B (based on DeepSeek). In this study, we selected Magicoder‑6.7B with the DeepSeek base model.
    DeepSeek-Coder Guo et al. (2024) is part of the DeepSeek family of code language models, developed by DeepSeek AI. The series includes pre-training and instruction-tuned models in parameter sizes from 1.3B up to 33B. We selected the 7B version for DeepSeek-Coder in this study.